An airline charges the following baggage fees: $25 for the first bag and $35 for the second. Suppose 54% of passengers have no checked luggage, 34% have only one piece of checked luggage and 12% have two pieces. We suppose a negligible portion of people check more than two bags.
) The average baggage-related revenue per passenger is: $
(please round to the nearest cent)
b) The standard deviation of baggage-related revenue is: $
(please round to the nearest cent)
c) About how much revenue should the airline expect for a flight of 120 passengers? $
(please round to the nearest dollar)

Answer by Boreal(15235)

x*p(x)
0*.54=0
.34*25=$8.50
.12*60=$7.20
The expected value is this sum or $15.70 per passenger
For 120 passengers, the expected value is 120*$15.70 or $1884
the variance is -15.70^2*0.54=133.10
plus (25-15.70)^2*0.34=29.41
plus (60-15.70)^2*0.12=235.50
These add to 398.01 $^2
sqrt of this is the sd or $19.95